FY 2011 AGS Polarized Proton Run Plan

October 26, 2010

Vincent Schoefer

 

 

Overview

·        RHIC Spin Run 11

o   TBD:  Polarized proton beam injected into Booster->AGS

o   TBD:  AtR available

o   1/1:  Injection into Blue

 

Baseline goal for extraction: 65% polarization at 1.5e11 protons.

 

Commissioning goals

·        Verify magnet survey impact on 6theta beta beat (Vincent)

·        Make horizontal tune jump quads operational (Woody)

·        Commission orbit harmonic feedback system and make operational (Leif)

·        Tune jump quad application

·        Test new CNI boards and motion control (Anatoli, Haixin)

·        AGS BBQ tune meter preliminary study (Kevin Mernick)

·        Matching to RHIC 9 MHz

·        Simultaneous commissioning of EBIS Au and U beams in Booster->AGS->AtR

 

 

Startup Schedule

·        TBD:  AGS walkthrough and SWEEP (start of controlled access)

·        TBD: Access controls, power supply testing and hi-pot

·        TBD: (Friday) MCR checkout of BtA/AGS, verify AGS user 4 setup for p^

·        TBD: (Monday) AGS heat run, Linac on

·        TBD: Install CNI boards, close up, reload p^ user functions (over heat run functions)

·        TBD: Injection into Booster

 

 

Beam Plans

·        Machine parameters:

o   ν_x=8.70, ν_y~8.95 at injection(archive from last run).

o   At injection, transverse emittance 10π (95% normalized); longitudinal emittance <0.8eV-s. >

o   Warm snake current=2670A.(8.5% at injection, 5.9% at extraction).

o   Cold snake helix current=238A and solenoid current=152A. (15.8% at injection, 10% at extraction).

o   AGS rep rate=4 sec.

o   The injection and extraction energies: = 4.5 and = 45.5, respectively.


Setup with Both Snakes Off

·        Establish beam in the Booster and then the AGS with both snakes off. [1-2 shifts]

·        Setup RF to establish beam at flattop. [1-2 shifts]

·        REV-TICK and Tunemeter setup. The tunemeter is critical to check tunes when snakes are on. As soon as it is done, we should move forward to snake on machine. [0-1 shift]

·        Move to HIGH VERTICAL TUNE at extraction, measure beta beat:  sin/cos response + midnight ORM

·        CNI polarimeter setup. The setup can start as soon as beam survives at extraction energy. It may take more than two shifts allocated here, but it can be stopped and continue after snake on. Prior to this work, the orbits (both horizontal and vertical) at top energy in C15 section should be set as flat (both in time and angle) as possible. Since the polarimeter setup involves different crew, it can be done in parallel with other activities. [1-2 shifts]

 

Setup with Both Snakes On

·        Establish beam in the AGS (injection at =4.5, extraction at = 45.5 with snake on. Start with on-demand mode and intensity of a few 10^10. Turn on cold snake (helical 238A, solenoid 152A,trim1 0A, trim2 7A) and E20 warm snake (2670A). Turn on correction quads and bumps for both snakes. While tuning the injection beam survival(first with rf off, then rf on), the beam dump bump (J10) should be on at injection porch. After reaching a few times 10^10 beam survives, remove the dump bump and tune for beam survival along the ramp. It may be efficient to load archives from last run and get beam circulating first with that setting. [3-5 shifts]

·        Commissioning instrumentation if not finished before: AGS tune-meter; IPM. [0-1 shift]

·        CNI polarimeter commissioning if not finished yet before. Since the polarimeter setup involves different crew, it can be done in parallel with other activities. [2 shifts].

·        Ramp development:

o   Tune the AGS injection to maximize beam survive at injection. Minimize emittances at AGS extraction (judged by IPM with rf off).

o   Correct both horizontal and vertical orbits with iterations.

o   Tune the 8th and 9th harmonics along the ramp.

o   Start with vertical tune around 8.8 along the ramp.

o   Raise vertical tunes along the ramp while tuning the orbit. The vertical tune should be chosen to avoid the higher order snake resonances, especially at 36+ν. The proposed vertical tune values are: 0+ν (~8.96 or higher), 12+ν (~8.97) , 36-ν(8.98) and 36+ν(8.985).

o   Measure polarization at extraction.
[4-6 shifts]

 

·        Measure polarization at extraction as function of B field. The field scan should cover = 44.5 to 46. [4 shifts]

·        Increase intensity from source gradually to the desired intensity. Measure polarization as intensity goes up. [2 shift]

·        Measure polarization at extraction as function of various quantities. There are several potential causes for polarization loss and they need to be scanned. The purpose of these scans are to find the optimum working point, so there is no need to make wide range scan. [4 shifts]

o   9th harmonics in the early part of the ramp.

o   Vertical tune not high enough before 0+ν.

o   Vertical tune not high enough at 0+ν (high order snake resonance).

o   Vertical tune not high enough at 36+ν (high order snake resonance).

 

·        Harmonics corrections in Booster for both =3 and 4, sine and cosine. This can be done at flattop energy with large current variation (step size of 10A). [1-2 shifts]

·        Expect to reach 50-65% polarization (with 80% as input) and 1-1.5*10^11 intensity. When this goal is reached, the remaining time will be divided into three parts: Tune jump quads test; Up and down ramp polarization measurements; CNI polarimeter test.

 

Tune Jump Quads

·  Measure betatron tune along the ramp with each of the two jump quads on and off. Setup the tune jump quads strength based on the tune measurements. Setup the precise tune jump timing based on the tune measurements. [4 shifts]

·  Setup the orbits centered at the two tune jump quads. Measure beam emittances to see improvement. It may requires several iterations. Commission ‘smoothing’ mode of orbit feedback.[1 shifts]

·  Repeated polarization measurements at =45.5 with jump quads on and off at beam center. [2 shifts]

·  If not seen polarization gain with quads on, take repeated polarization measurements with jump quads on and off at beam center at =45.5 with tunemeter kick on. [2 shifts]

·  Repeated polarization measurements at =45.5 with jump quads timing shifted ( timing scan). [4 shifts]

·  After seen the jump quads effect, take horizontal polarization profiles at =45.5 with jump quadson and off. [2 shifts]

 

EBIS

Work in the Booster with He and Ne likely to occupy EBIS until Thanksgiving (11/24).  Thereafter, work in the Booster with Au (and maybe U) will commence (PPM with p^ from Linac). Time-sharing of AGS thereafter contingent on polarization progress.

 

AGS pp Operation System Experts


Booster scraping: Keith
BtA lattice: Keith, Nick, Kevin
AGS/Booster optics control application: Kip, Kevin
AGS injection tuning: Keith, Leif, Woody
AGS RF: Keith, KevinS
AGS tune-meter: Leif
AGS IPM: Kevin
AGS orbit control and display: Leif, Woody
Data spreadsheet: Seth
AGS snake operation: Haixin, Leif
Snake bumps and corrections: Nick, Leif, Alfredo
AGS E880 polarimeter: Haixin
AGS CNI polarimeter: Kin, Haixin
AGS extraction: Vincent, Greg, Nick
AGS lattice matching: Alfredo, Nick